University of Illinois Department of Economics Spring 2001 Roger Koenker

Economics 476 Term Paper Assignment

“You haven’t told me yet,” said Lady Nuttal, ”What it is your fiance does for a living.”

“He’s a statistician,” replied Lamia, with an annoying sense of being on the defensive.

Lady Nuttal was obviously taken aback. It had not occurred to her that statisticians entered into normal social relationships. The species, she would have surmised, was perpetuated in some collateral manner, like mules.

“But Aunt Sara, it’s a very interesting profession,” said Lamia warmly.

“I don’t doubt it,” said her aunt, who obviously doubted it very much. “To express anything important in mere figures is so plainly impossible that there must be endless scope for well-paid advice on how to do it. But don’t you think that life with a statistician would be rather, shall we say, humdrum?”

Lamia was silent. She felt reluctant to discuss the surprising depth of emotional possibility which she had discovered below Edward’s numerical veneer.

“It’s not the figures themselves,” she said finally, “It’s what you do with them that matters.”

K.A.C. Mandeville, The Undoing of Lamia Gurdleneck

A brief (5-15 pages) term paper will be expected from each student by Noon, May 4. The paper may take one of three forms:

• a theoretical note on some aspect of econometrics we have touched upon in class, an elaboration of one of the problems for example.

• an empirical note which explores the application of a technique we have considered on a “real” data set.

• a critical survey of recent developments in a relevant niche of the econometrics literature, the emphasis here should be on the word critical. You should try to offer some guidance on where the literature should go from here.

While I don’t encourage purely “monte carlo” papers, some monte-carlo aspect is encouraged. To avoid potential misunderstandings about paper topics, I would like all students to submit a 1-page “paper proposal by March 9. This is the last day of class before spring break. These proposals should contain at least three relevant references to literature and a brief sketch of what you propose to do in your paper.